# Bidirectional Gated Mamba for Sequential Recommendation

### v2

@@
Abstract.
In various domains, Sequential Recommender Systems (SRS) have become essential due to their superior capability to discern intricate user preferences. Typically, SRS utilize transformer-based architectures to forecast the subsequent item within a sequence. Nevertheless, the quadratic computational complexity inherent in these models often leads to inefficiencies, hindering the achievement of real-time recommendations. Mamba, a recent advancement, has exhibited exceptional performance in time series prediction, significantly enhancing both efficiency and accuracy. However, integrating Mamba directly into SRS poses several challenges. Its inherently unidirectional nature may constrain the model’s capacity to capture the full context of user-item interactions, while its instability in state estimation can compromise its ability to detect short-term patterns within interaction sequences.
To overcome these issues, we introduce a new framework named SelectIve Gated MAmba (SIGMA). This framework leverages a Partially Flipped Mamba (PF-Mamba) to construct a bidirectional architecture specifically tailored to improve contextual modeling. Additionally, an input-sensitive Dense Selective Gate (DS Gate) is employed to optimize directional weights and enhance the processing of sequential information in PF-Mamba. For short sequence modeling, we have also developed a Feature Extract GRU (FE-GRU) to efficiently capture short-term dependencies. Empirical results indicate that SIGMA outperforms current models on five real-world datasets. Our implementation code is available at https://github.com/ziwliu-cityu/SIMGA to ease reproducibility.
Sequential Recommender Systems, State Space Model, Gate Mechanism
journalyear: 2024ccs: Information systems Recommender systems
1. Introduction
@@
Over the past decade, sequential recommender systems (SRS) have demonstrated promising potential across various domains, including content streaming platforms (Rappaz et al., 2021) and e-commerce (Wang et al., 2020). To harness this potential and meet the demand for accurate next-item predictions (Fang et al., 2020), an increasing number of researchers are focusing on refining existing architectures and proposing novel approaches (Wang et al., 2019; Miller et al., 2024).
Figure 1. The illustration for long-tail user problem.
@@
Recently, Transformer-based models have emerged as the leading approach in sequential recommendation due to their outstanding performance (de Souza Pereira Moreira et al., 2021). By leveraging the powerful self-attention mechanism (Vaswani et al., 2017; Keles et al., 2023), these models have demonstrated a remarkable ability to deliver accurate predictions. However, despite their impressive performance, current transformer-based models are proven inefficient since the amount of computation grows quadratically as the length of the input sequence increases (Keles et al., 2023). Other approaches, such as RNN-based models e.g., GRU4Rec (Jannach and Ludewig, 2017) and MLP-based models e.g., MLP4Rec (Li et al., 2022a), are proven to be efficient due to their linear complexity. Nevertheless, they have struggled with handling long and complex patterns (Yoon and Jang, 2023). All these methods above seem to have suffered from a significant trade-off between effectiveness and efficiency. Consequently, a specially designed State Space Model (SSM) called Mamba (Gu and Dao, 2023) has been proposed. By employing simple input-dependent selection on the original SSM (Liu et al., 2024a; Hamilton, 1994), it has demonstrated remarkable efficiency and effectiveness.
@@
Figure 2. Framework of proposed SIGMA. The core part of this framework is the G-Mamba Block, which can directly tackle the context modeling and short sequence modeling challenges when introducing Mamba to SRS.
@@
  • Context Modeling: While previous research has demonstrated Mamba’s reliability in capturing sequential information (Gu and Dao, 2023; Yang et al., 2024), its unidirectional architecture imposes significant limitations in SRS applications. By only capturing users’ past behaviors, Mamba cannot leverage future contextual information, potentially leading to an incomplete understanding of user preferences  (Liu et al., 2024a; Sun et al., 2019). For instance, if a user consistently purchases household items but begins to show interest in sports equipment, a model that does not consider future contexts may struggle to recognize this shift, resulting in sub-optimal next-item predictions (Jiang et al., 2024; Kweon et al., 2021).
  • Short Sequence Modeling: This challenge is primarily driven by the long-tail user problem, a common issue in sequential recommendation. Long-tail users refer to such users who interact with only a few items but typically receive lower-quality recommendations compared to the normal ones (Kim et al., 2019a, b). Furthermore, the instability in state estimation caused by limited data in short sequences (Gu and Dao, 2023; Smith et al., 2022; Yu et al., 2024) exacerbates this problem when Mamba is directly applied to SRS, highlighting the need for effective short sequence modeling. For illustration, we compare two leading baselines, Mamba4Rec (Liu et al., 2024a) and SASRec (Kang and McAuley, 2018), against our proposed framework on the Beauty dataset. As shown in Figure 1, the histogram depicts the number of users in each group, while the line represents recommendation performance in terms of Hit@10. SASRec outperforms Mamba4Rec in the first three groups, demonstrating Mamba4Rec’s exacerbation of the long-tail user problem.
To address these challenges and better leverage Mamba’s strengths, we propose an innovative framework called SelectIve Gated MAmba for Sequential Recommendation (SIGMA). Our approach introduces the Partially Flipped Mamba (PF-Mamba), a specialized bidirectional structure that captures contextual information (Liu et al., 2024a; Jiang et al., 2024). We then introduce an input-dependent Dense Selective Gate (DS Gate) to allocate the weights of the two directions and further filter the information. Additionally, we develop a Feature Extract GRU (FE-GRU) to better model short-term patterns in interaction sequences (Hidasi et al., 2015), offering a possible solution to the long-tail user problem. Our contributions are summarized as follows::
@@
2. Methodology
In this section, we will introduce a novel framework, SIGMA, which effectively addresses the aforementioned problems by adopting PF-Mamba with a Dense Selective Gate and a Feature Extract GRU. We will first present the overview of our proposed framework; then detail the important components of our architecture; and lastly introduce how we conduct our training and inference procedures.
2.1. Framework Overview
In this section, we present an overview of our proposed framework in Figure 2.
Firstly, we employ an embedding layer to learn the representation for input items. After getting the high-dimensional interaction representation, we propose a G-Mamba block to selectively extract the information. Specifically, the G-Mamba block consists of a bidirectional Mamba path and a GRU path, which respectively address challenges in context modeling and short sequence modeling. Then, a Position-wise Feed-Forward Network (PFFN) is adopted to improve the modeling ability of user actions in the hidden representation. Finally, processed by the prediction layer, we can get an accurate next-item prediction.
2.2. Embedding Layer
For existing SRSs, It is necessary to map the sequential information in user-item interaction to a high-dimensional space (Liu et al., 2023a) to effectively capture the temporal dependencies. In our framework, we choose a commonly used method for constructing the item embedding. Here, we denote the user set as $\mathcal{U}=\left\{{u}_{1},{u}_{2},\cdots,{u}_{\left|\mathcal{U}\right|}\right\}$ and the item set as $\mathcal{V}=\left\{{v}_{1},{v}_{2},\cdots,{v}_{\left|\mathcal{V}\right|}\right\}$. So for a chronologically ordered interaction sequence, it can be expressed as $\boldsymbol{S}_{u}=\left[{s}_{1},{s}_{2},\cdots,{s}_{{n}_{u}}\right]$, where ${n}_{u}$ represents the length of the sequence for user $u\in\mathcal{U}$. For simplicity, we omit the mark $\left(u\right)$ in the following sections. Regarding this interaction sequence as the input tensor, we denote $D$ as the embedding dimension and use a learnable item embedding matrix $\boldsymbol{E}\in\mathbb{{R}^{\mathit{\left|\mathcal{V}\right|\times D}}}$ to adaptively projected ${s}_{i}$ into the representation $\boldsymbol{h}_{i}$. The whole interaction sequence can be output as:
(1) $\boldsymbol{H}_{0}=\left[\boldsymbol{{h}}_{1},\boldsymbol{h}_{2},\cdots,\boldsymbol{h}_{N}\right]$
@@ 2.2. Embedding Layer
2.3. G-Mamba Block
In this section, we will detail the design of our proposed G-Mamba Block. Starting with the input sequence processed by the Embedding Layer, this block introduces two paralleling paths i.e., PF-Mamba and FE-GRU, which respectively address the context modeling challenge and short sequence modeling challenge. Specifically, for the contextual information loss caused by the unidirectional structure of Mamba (Gu and Dao, 2023; Kweon et al., 2021), we introduce the Partially Flipped Mamba. It modifies the original unidirectional structure to a bi-directional one by employing a reverse block that partially flips the first $r$ items in the interaction sequence. Next, a Dense Selective Gate is proposed to properly allocate the weights of the two directions depending on the input sequence (Qin et al., 2024; Zhang et al., 2024). Additionally, for the long-tail user problem, we introduce the Feature Extract GRU, effectively capturing short-term preferences (Hidasi et al., 2015; Kim et al., 2019b).
2.3.1. Partially Flipped Mamba
This module is proposed to address the context modeling challenge by leveraging the bi-directional structure. Current bi-directional methods like Dual-path Mamba (Jiang et al., 2024) or Vision Mamba (Zhu et al., 2024) usually just flip the whole input sentence to enable the global capturing capability. Although it allows the model to have a better understanding of the context, it significantly reduces the influence of short-term patterns in interaction sequences, leading to the loss of important interest dependencies. To address this issue, we introduce a partial flip method and integrate it with the Mamba block to construct a bi-directional structure. Followed by embedding sequence $\boldsymbol{H}_{0}$ in Equation (1), the partially flip function adaptively reverses the first $n$ items while remaining the last $r$ items in the input tensor from $\boldsymbol{H}_{0}=\left[\boldsymbol{h}_{1},\boldsymbol{h}_{2},\cdots,\boldsymbol{h}_{n},\boldsymbol{h}_{n+1}\cdots,\boldsymbol{h}_{N}\right]$ to $\boldsymbol{H}_{0}^{f}=\left[\boldsymbol{h}_{n},\cdots,\boldsymbol{h}_{2},\boldsymbol{h}_{1},\boldsymbol{h}_{n+1},\cdots,\boldsymbol{h}_{N}\right]$. $r$ is a pre-defined hyperparameter that equals $N-n$, which determines the range of the remaining items, i.e., what extent we focus on the short-term preferences. After processing the input sequence, we utilize two Mamba blocks to construct a bi-directional architecture and process these two sequences as follows:
(2) $\boldsymbol{{M}}_{0}$ $=\operatorname{Mamba}\left(\boldsymbol{H}_{0}\right)\in\mathbb{R}^{L\times D}$
@@ 2.3.1. Partially Flipped Mamba
where ${L}$ and ${D}$ respectively represent the sequence length and hidden dimension at each time step. These two feature representations will then get a dot product with an input-dependent DS Gate to further learn the user preferences.
(3) $\hat{\boldsymbol{M}}_{0}=\mathcal{G}_{1}\left(\boldsymbol{H}_{0}\right)\cdot\boldsymbol{M}_{0}+\mathcal{G}_{1}\left(\boldsymbol{H}_{0}^{f}\right)\cdot\boldsymbol{M}_{0}^{f}$
where $\mathcal{G}_{1}$ represents the designed DS Gate and $\hat{\boldsymbol{M}}_{0}={\left[\boldsymbol{m}_{1},\boldsymbol{m}_{2},\cdots,\boldsymbol{m}_{N}\right]}^{T}$ denotes the output from PF-Mamba.
2.3.2. Dense Selective Gate
To allocate the weights of two Mamba blocks and further filter the information according to the input sequence, we design an input-dependent Dense Selective Gate. It starts with a dense layer and a Conv1d layer to extract the original sequential information from the context, which can be formalized as follows:
(4) $\boldsymbol{G}_{0}=\operatorname{Conv1d}\left(\boldsymbol{H}_{0}\boldsymbol{W}_{\sigma}^{\mathit{(1)}}+b_{\sigma}^{\mathit{(1)}}\right)$
where $\boldsymbol{H}_{0}$ is denoted as the output of embedding layer followed by Equation (1). Then, we introduce a forget gate and a $\operatorname{SiLU}$ gate (Qin et al., 2024) to generate the weights from the interaction sequence:
(5) $\boldsymbol{\delta}_{1}\left(\boldsymbol{{G}}_{0}\right)$ $=\boldsymbol{{G}}_{0}\boldsymbol{W}_{\delta}^{(1)}+b_{\delta}^{(1)}$
@@ 2.3.2. Dense Selective Gate
where $W_{\delta}^{(1)}\in\mathbb{R}^{D\times D}$ is the weight, $\boldsymbol{b}_{\delta}^{(1)}\in\mathbb{R}^{D}$ is bias; $\mathcal{G}_{0}$ denoted as the symbol of forget gate; $\sigma(\cdot)$ represents the Sigmoid activation function (He et al., 2018). By employing this $\mathcal{G}_{0}$, we can control the information flow in $\boldsymbol{G}_{0}$ to selectively retain or suppress certain information (De et al., 2024). Apart from the $\mathcal{G}_{0}$, we also employ a $\operatorname{SiLU}$ function to further improve the capability for capturing more complex patterns and features (Nwankpa et al., 2018). Therefore, We can conclude our DS Gate as follows:
(6) $\mathcal{G}_{1}(\boldsymbol{H}_{0})=\operatorname{SiLU}\left(\boldsymbol{\delta}_{1}(\boldsymbol{G}_{0})\right)+\mathcal{G}_{0}(\boldsymbol{G}_{0})$
@@
2.3.3. Feature Extract GRU
To handle Mamba’s undesirable performance on short sequence modeling, we introduce one more GRU path called Feature Extract GRU in our SIGMA framework. Considering efficiency and effectiveness, we only introduce one more convolution function before the GRU cell to extract and mix the features (Yuan et al., 2019). By employing this one-dimensional convolution with a well-designed kernel size, we can aggregate and extract information from the short-term pattern of the input embedding sequence. Then, we can extract the hidden representation by utilizing GRU’s impressive capability in capturing short-term dependencies. The whole processing procedure can then be formalized as follows:
@@ 2.3.3. Feature Extract GRU
(8) $\boldsymbol{Z}_{0}={a}_{1}\boldsymbol{{M}}+{a}_{2}\boldsymbol{F}_{0}\in\mathbb{R}^{\mathit{L\times D}}$
@@
(9) $\hat{\boldsymbol{Z}}_{0}=\boldsymbol{{Z}}_{0}\boldsymbol{W}_{\delta}^{(2)}+\boldsymbol{b}_{\delta}^{(2)}$
@@
2.4. PFFN Network
To capture the complex features, we further leverage a position-wise feed-forward network (PFFN Net) (Liu et al., 2024a; Kang and McAuley, 2018):
(10) $\boldsymbol{R}_{0}=\operatorname{GELU}\left(\hat{\boldsymbol{Z}}_{0}\boldsymbol{W}_{\delta}^{(3)}+b_{\delta}^{(3)}\right)\boldsymbol{W}_{\delta}^{(4)}+b_{\delta}^{(4)}$
@@ 2.5. Train and Inference
(11) $\text{logits}_{ik}=\sum_{j=1}^{d}\mathbf{R}_{ij}\cdot\mathbf{H}_{kj}^{item}$
@@
Where $\text{logits}_{ik}$ and $\text{P}_{ik}$ respectively represent the prediction scores and corresponding probability of the i-th sample for the k-th item. And we can formulate our Cross Entropy Loss (CE) (Zhang and Sabuncu, 2018) and minimize it as:
(12) $\mathcal{L}_{CE}=-\frac{1}{B}\sum_{i=1}^{B}\log P_{i,y_{i}}$
Where ${y}_{i}$ represents the actual positive sample for i-th sample and ${B}$ represents the batch size following the previous formulas. By constantly updating the loss in each epoch, we can obtain the optimal weighting parameters and correspondingly get an accurate next-item prediction.
Table 1. The Statistics of Datasets
@@
Table 2. Overall performance comparison between SIGMA and other baselines. The best results are bold, and the second-best are underlined. “*” indicates the improvements are statistically significant (i.e., two-sided t-test with $p<0.05$ ) over baselines).

### v3

# SIGMA: Selective Gated Mamba for Sequential Recommendation

@@
Abstract
Sequential Recommender Systems (SRS) have emerged as a promising technique across various domains, excelling at capturing complex user preferences. Current SRS have employed transformer-based models to give the next-item prediction. However, their quadratic computational complexity often lead to notable inefficiencies, posing a significant obstacle to real-time recommendation processes. Recently, Mamba has demonstrated its exceptional effectiveness in time series prediction, delivering substantial improvements in both efficiency and effectiveness. However, directly applying Mamba to SRS poses certain challenges. Its unidirectional structure may impede the ability to capture contextual information in user-item interactions, while its instability in state estimation may hinder the ability to capture short-term patterns in interaction sequences. To address these issues, we propose a novel framework called SelectIve Gated MAmba for Sequential Recommendation (SIGMA). By introducing the Partially Flipped Mamba (PF-Mamba), we construct a special bi-directional structure to address the context modeling challenge. Then, to consolidate PF-Mamba’s performance, we employ an input-dependent Dense Selective Gate (DS Gate) to allocate the weights of the two directions and further filter the sequential information. Moreover, for short sequence modeling, we devise a Feature Extract GRU (FE-GRU) to capture the short-term dependencies. Experimental results demonstrate that SIGMA significantly outperforms existing baselines across five real-world datasets. Our implementation code is available at https://github.com/Applied-Machine-Learning-Lab/SIMGA.
Introduction
@@
Over the past decade, sequential recommender systems (SRS) have demonstrated promising potential across various domains, including content streaming platforms (Song et al. 2022; Zhao et al. 2023a), e-commerce (Wang et al. 2020) and other domains (Li et al. 2022). To harness this potential and meet the demand for accurate next-item predictions (Fang et al. 2020; Liu et al. 2023c), an increasing number of researchers are focusing on refining existing architectures and proposing novel approaches (Wang et al. 2019; Liu et al. 2024b; Wang et al. 2023).
Figure 1: The illustration for long-tail user problem.
@@
Recently, Transformer-based models have emerged as the leading approaches in sequential recommendation due to their outstanding performance (de Souza Pereira Moreira et al. 2021). By leveraging the powerful self-attention mechanism (Vaswani et al. 2017; Keles, Wijewardena, and Hegde 2023), these models have demonstrated a remarkable ability to deliver accurate predictions. However, despite their impressive performance, current transformer-based models are proven inefficient since the amount of computation grows quadratically as the length of the input sequence increases (Keles, Wijewardena, and Hegde 2023). Other approaches, such as RNN-based models (Jannach and Ludewig 2017) and MLP-based models (Li et al. 2023b; Gao et al. 2024; Liang et al. 2023), are proven to be efficient due to their linear complexity. Nevertheless, they have struggled with handling long and complex patterns (Yoon and Jang 2023). All these methods above seem to have suffered from a significant trade-off between effectiveness and efficiency. Consequently, a specially designed State Space Model (SSM) called Mamba (Gu and Dao 2023) has been proposed. By employing simple input-dependent selection on the original SSM (Liu et al. 2024a; Hamilton 1994), it has demonstrated remarkable efficiency and effectiveness.
@@
Figure 2: Framework of proposed SIGMA. The core part of this framework is the G-Mamba Block, which can directly tackle the context modeling and short sequence modeling challenges when introducing Mamba to SRS.
@@
  • Context Modeling: While previous researches has demonstrated Mamba’s reliability in capturing sequential information (Gu and Dao 2023; Yang et al. 2024), its unidirectional architecture imposes significant limitations when applying to SRS. By only capturing users’ past behaviors, Mamba can not leverage future contextual information, potentially leading to an incomplete understanding of users’ preferences  (Liu et al. 2024a; Sun et al. 2019). For instance, if a user consistently purchases household items but begins to show interest in sports equipment, a model that does not consider future context may struggle to recognize this shift, resulting in sub-optimal next-item predictions (Jiang, Han, and Mesgarani 2024; Kweon, Kang, and Yu 2021).
  • Short Sequence Modeling: This challenge is primarily driven by the long-tail user problem, a common issue in sequential recommendation. Long-tail users refer to such users who interact with only a few items but typically receive lower-quality recommendations compared to the normal ones (Kim et al. 2019a, b; Liu et al. 2024d). Furthermore, the instability in state estimation caused by limited data in short sequences (Gu and Dao 2023; Smith, Warrington, and Linderman 2022; Yu, Mahoney, and Erichson 2024) exacerbates this problem when Mamba is directly applied to SRS, highlighting the need for effectively modeling short sequences. For illustration, we compare two leading baselines, Mamba4Rec (Liu et al. 2024a) and SASRec (Kang and McAuley 2018), against our proposed framework on the Beauty dataset. As shown in Figure 1, the histogram depicts the number of users in each group, while the line represents recommendation performance in terms of Hit@10. SASRec outperforms Mamba4Rec in the first three groups, demonstrating Mamba4Rec’s exacerbation of the long-tail user problem.
To address these challenges and better leverage Mamba’s strengths, we propose an innovative framework called SelectIve Gated MAmba for Sequential Recommendation (SIGMA). Our approach introduces the Partially Flipped Mamba (PF-Mamba), a specialized bidirectional structure that captures contextual information (Liu et al. 2024a; Jiang, Han, and Mesgarani 2024). We then introduce an input-dependent Dense Selective Gate (DS Gate) to allocate the weights of the two directions and further filter the information. Additionally, we develop a Feature Extract GRU (FE-GRU) to better model short-term patterns in interaction sequences (Hidasi et al. 2015), offering a possible solution to the long-tail user problem. Our contributions are summarized as follows:
@@
Methodology
In this section, we will introduce a novel framework, SIGMA, which effectively addresses the aforementioned problems by adopting PF-Mamba with a Dense Selective Gate and a Feature Extract GRU. We will first present an overview of our proposed framework; then detail the important components of our architecture; and lastly introduce how we conduct our training and inference procedures.
Framework Overview
In this section, we present an overview of our proposed framework in Figure 2. Firstly, we employ an embedding layer to learn the representation for input items. After getting the high-dimensional interaction representation, we propose a G-Mamba block to selectively extract the information. Specifically, the G-Mamba block consists of a bidirectional Mamba path and a GRU path, which respectively address challenges in context modeling and short sequence modeling. Then, a Position-wise Feed-Forward Network (PFFN) is adopted to improve the modeling ability of users’ actions in the hidden representation. Finally, processed by the prediction layer, we can get the accurate next-item predictions.
Embedding Layer
For existing SRS, It is necessary to map the sequential information in user-item interaction to a high-dimensional space (Zhao et al. 2023b) to effectively capture the temporal dependencies. In our framework, we choose a commonly used method for constructing the item embedding. Here, we denote the user set as $\mathcal{U}=\left\{{u}_{1},{u}_{2},\cdots,{u}_{\left|\mathcal{U}\right|}\right\}$ and the item set as $\mathcal{V}=\left\{{v}_{1},{v}_{2},\cdots,{v}_{\left|\mathcal{V}\right|}\right\}$. So for a chronologically ordered interaction sequence, it can be expressed as $\boldsymbol{S}_{u}=\left[{s}_{1},{s}_{2},\cdots,{s}_{{n}_{u}}\right]$, where ${n}_{u}$ represents the length of the sequence for user $u\in\mathcal{U}$. For simplicity, we omit the mark $\left(u\right)$ in the following sections. Regarding this interaction sequence as the input tensor, we denote $D$ as the embedding dimension and use a learnable item embedding matrix $\boldsymbol{E}\in\mathbb{{R}^{\mathit{\left|\mathcal{V}\right|\times D}}}$ to adaptively projected ${s}_{i}$ into the representation $\boldsymbol{h}_{i}$. The whole interaction sequence can be output as:
$\boldsymbol{H}_{0}=\left[\boldsymbol{{h}}_{1},\boldsymbol{h}_{2},\cdots,\boldsymbol{h}_{N}\right]$ (1)
@@
G-Mamba Block
In this section, we will detail the design of our proposed G-Mamba Block. Starting with the input sequence processed by the Embedding Layer, this block introduces two paralleling paths i.e., PF-Mamba and FE-GRU, which respectively address the context modeling challenge and short sequence modeling challenge. Specifically, for the contextual information loss caused by the unidirectional structure of Mamba (Gu and Dao 2023; Kweon, Kang, and Yu 2021), we introduce the Partially Flipped Mamba. It modifies the original unidirectional structure to a bi-directional one by employing a reverse block that retains the last r items while flipping the preceding items. Next, a Dense Selective Gate is proposed to properly allocate the weights of the two directions depending on the input sequence (Qin, Yang, and Zhong 2024; Zhang, Wang, and Zhao 2024). Additionally, for the long-tail user problem, we introduce the Feature Extract GRU to capture short-term preferences effectively (Hidasi et al. 2015; Kim et al. 2019b).
Partially Flipped Mamba. This module is proposed to address the context modeling challenge by leveraging the bi-directional structure. Current bi-directional methods like Dual-path Mamba (Jiang, Han, and Mesgarani 2024) or Vision Mamba (Zhu et al. 2024) usually just flip the whole input sentence to enable the global capturing capability. Although it allows the model to have a better understanding of the context, it significantly reduces the influence of short-term patterns in interaction sequences, leading to the loss of important interest dependencies. To address this issue, we introduce a partial flip method and integrate it with the Mamba block to construct a bi-directional structure. Followed by embedding sequence $\boldsymbol{H}_{0}$ in Equation (1), the partially flip function adaptively reverses the first $n$ items while remaining the last $r$ items in the input tensor from $\boldsymbol{H}_{0}=\left[\boldsymbol{h}_{1},\boldsymbol{h}_{2},\cdots,\boldsymbol{h}_{n},\boldsymbol{h}_{n+1}\cdots,\boldsymbol{h}_{N}\right]$ to $\boldsymbol{H}_{0}^{f}=\left[\boldsymbol{h}_{n},\cdots,\boldsymbol{h}_{2},\boldsymbol{h}_{1},\boldsymbol{h}_{n+1},\cdots,\boldsymbol{h}_{N}\right]$. $r$ is a pre-defined hyperparameter that equals $N-n$, which determines the range of the remaining items, i.e., what extent we focus on the short-term preferences. After processing the input sequence, we utilize two Mamba blocks to construct a bi-directional architecture and process these two sequences as follows:
$\boldsymbol{{M}}_{0}$ $=\operatorname{Mamba}\left(\boldsymbol{H}_{0}\right)\in\mathbb{R}^{L\times D}$ (2)
@@
where ${L}$ and ${D}$ respectively represent the sequence length and hidden dimension. These two feature representations will then get a dot product with an input-dependent DS Gate to further learn the user preferences.
$\hat{\boldsymbol{M}}_{0}=\mathcal{G}_{1}\left(\boldsymbol{H}_{0}\right)\cdot\boldsymbol{M}_{0}+\mathcal{G}_{1}\left(\boldsymbol{H}_{0}^{f}\right)\cdot\boldsymbol{M}_{0}^{f}$ (3)
where $\mathcal{G}_{1}$ represents the designed DS Gate and $\hat{\boldsymbol{M}}_{0}={\left[\boldsymbol{m}_{1},\boldsymbol{m}_{2},\cdots,\boldsymbol{m}_{N}\right]}^{T}$ denotes the output from PF-Mamba. Dense Selective Gate. To allocate the weights of two Mamba blocks and further filter the information according to the input sequence, we design an input-dependent Dense Selective Gate. It starts with a dense layer and a Conv1d layer to extract the original sequential information from the context, which can be formalized as follows:
$\boldsymbol{G}_{0}=\operatorname{Conv1d}\left(\boldsymbol{H}_{0}\boldsymbol{W}_{\sigma}^{\mathit{(1)}}+b_{\sigma}^{\mathit{(1)}}\right)$ (4)
where $\boldsymbol{H}_{0}$ is denoted as the output of embedding layer followed by Equation (1). Then, we introduce a forget gate and a $\operatorname{SiLU}$ gate (Qin, Yang, and Zhong 2024) to generate the weights from the interaction sequence:
$\boldsymbol{\delta}_{1}\left(\boldsymbol{{G}}_{0}\right)$ $=\boldsymbol{{G}}_{0}\boldsymbol{W}_{\delta}^{(1)}+b_{\delta}^{(1)}$ (5)
@@
where $W_{\delta}^{(1)}\in\mathbb{R}^{D\times D}$ is the weight, $\boldsymbol{b}_{\delta}^{(1)}\in\mathbb{R}^{D}$ is bias; $\mathcal{G}_{0}$ is denoted as the symbol of forget gate; $\sigma(\cdot)$ represents the Sigmoid activation function (He et al. 2018). By employing this $\mathcal{G}_{0}$, we can control the information flow in $\boldsymbol{G}_{0}$ to selectively retain or suppress certain information (De et al. 2024). Apart from the $\mathcal{G}_{0}$, we also employ a $\operatorname{SiLU}$ function to further improve the capability for capturing more complex patterns and features (Nwankpa et al. 2018). Therefore, We can conclude our DS Gate as follows:
$\mathcal{G}_{1}(\boldsymbol{H}_{0})=\operatorname{SiLU}\left(\boldsymbol{\delta}_{1}(\boldsymbol{G}_{0})\right)+\mathcal{G}_{0}(\boldsymbol{G}_{0})$ (6)
@@
Feature Extract GRU. To handle Mamba’s undesirable performance on short sequence modeling, we introduce one more GRU path called Feature Extract GRU in our SIGMA framework. Considering efficiency and effectiveness, we only introduce one convolution function before the GRU cell to extract and mix the features (Yuan et al. 2019). By employing this one-dimensional convolution with a well-designed kernel size, we can aggregate and extract information from the short-term pattern of the input embedding sequence. Then, we can extract the hidden representation by utilizing GRU’s impressive capability to capture short-term dependencies. The whole processing procedure can then be formalized as follows:
@@
$\boldsymbol{Z}_{0}={a}_{1}\boldsymbol{{M}}+{a}_{2}\boldsymbol{F}_{0}\in\mathbb{R}^{\mathit{L\times D}}$ (8)
@@
$\hat{\boldsymbol{Z}}_{0}=\boldsymbol{{Z}}_{0}\boldsymbol{W}_{\delta}^{(2)}+\boldsymbol{b}_{\delta}^{(2)}$ (9)
@@
PFFN Network
To capture the complex features, we further leverage a position-wise feed-forward network (PFFN Net) (Liu et al. 2024a; Kang and McAuley 2018):
$\boldsymbol{R}_{0}=\operatorname{GELU}\left(\hat{\boldsymbol{Z}}_{0}\boldsymbol{W}_{\delta}^{(3)}+b_{\delta}^{(3)}\right)\boldsymbol{W}_{\delta}^{(4)}+b_{\delta}^{(4)}$ (10)
@@
$\text{logits}_{ik}=\sum_{j=1}^{d}\mathbf{R}_{ij}\cdot\mathbf{H}_{kj}^{item}$ (11)
@@
Where $\text{logits}_{ik}$ and $\text{P}_{ik}$ respectively represent the prediction score and corresponding probability of the i-th sample for the k-th item. Correspondingly, we can formulate our Cross Entropy Loss (CE) (Zhang and Sabuncu 2018) and minimize it as:
$\mathcal{L}_{CE}=-\frac{1}{B}\sum_{i=1}^{B}\log P_{i,y_{i}}$ (12)
Where ${y}_{i}$ represents the actual positive sample for i-th sample and ${B}$ represents the batch size. By constantly updating the loss in each epoch, we can obtain the optimal weighting parameters and correspondingly get an accurate next-item prediction.
@@
Table 1: The statistics of datasets